OpenStudy (anonymous):

Write an inequality 3y=|x-3| I know how to write inequalities but I don't get what I am supposed to do with this one. I got y=1/3x-1 but I have a feeling I did it wrong. If I did could someone show me steps one how to do it? Thanks.

OpenStudy (anonymous):

leave equation as is and break it into two parts like this \[3y = x - 3\] and \[-3y = x - 3\] now solve for y in each of the above equation. then set the equations equal to each other and find the value of x. once you find the value of x, plug it into the original equation and ask yourself, "what inequality do I need to use to make it true?"